OpenStudy (anonymous):

Solve the following system of equations. Please show your work to receive full credit. x - y = 19 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

-2(x-y=19) 2x+y=2 -2x+2y=-38 2x+ y= 2 ___________ 3y=-36 Y= -12 Then plug in y for on of the equations 2x+(-12)=2 2x=14 x=7 So you used the elimination method

OpenStudy (anonymous):

thanks :)